OpenStudy (howard-wolowitz):

Ok, I got this : 4.0460502 but am I correct?

OpenStudy (howard-wolowitz):

The answer I got is D!

Nnesha (nnesha):

and your work ?? i'm sorry but tht's important

OpenStudy (anonymous):

yeah good. 4.04 something..

OpenStudy (howard-wolowitz):

dont need my work because I"m correct :)

Nnesha (nnesha):

but still i would like to see it

OpenStudy (howard-wolowitz):

unless you just wana see it

Nnesha (nnesha):

yes i want to see it

OpenStudy (anonymous):

You need to show work if you are wrong, as to figure out where you have done wrong, but if you are correct, then you are correct, no need to show your work..

OpenStudy (howard-wolowitz):

Excatly, and I"m right BIG GURL

Nnesha (nnesha):

nope still work is important

OpenStudy (howard-wolowitz):

*exactly

OpenStudy (anonymous):

And when @Nnesha wants to see your work, then show her, may be finally you come out to be wrong.. :P

OpenStudy (howard-wolowitz):

ok fine here:

OpenStudy (howard-wolowitz):

ok fine here: Remove parentheses around 0.65. 4000⋅0.65z=700 Divide each term in the equation by 4000.

OpenStudy (howard-wolowitz):

then : 0.65^z=740 Divide each expression in the equation by 4000. 0.65^z=7004000

OpenStudy (anonymous):

What are you doing man.. You are insulting me.. :P

Nnesha (nnesha):

^

OpenStudy (howard-wolowitz):

then : 0.65^z=740 Take the natural logarithm of both sides of the equation to remove the variable from the exponent. ln(0.65^z)=ln(740)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Howard-Wolowitz ok fine here: Remove parentheses around 0.65. 4000⋅0.65z=700 Divide each term in the equation by 4000. \(\color{blue}{\text{End of Quote}}\) here we go1!!!!

OpenStudy (howard-wolowitz):

what do you mean?

OpenStudy (howard-wolowitz):

wont me to stop

OpenStudy (anonymous):

How you get 740 on right side?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Howard-Wolowitz ok fine here: Remove parentheses around 0.65. 4000⋅0.65z=700 Divide each term in the equation by 4000. \(\color{blue}{\text{End of Quote}}\) you can't just remove the parentheses :(

OpenStudy (howard-wolowitz):

@radar am I correct with D

OpenStudy (howard-wolowitz):

its not 740 its 7/40

OpenStudy (howard-wolowitz):

The logarithm of a product is equal to the sum of the logarithms of each factor (e.g.log(xy)=log(x)+log(y)). The logarithm of a division is equal to the difference of the logarithms of each factor (e.g.log(xy)=log(x)−log(y)). −0.43078292z=1(ln(7)−ln(40))

Nnesha (nnesha):

okay do mean (.65)^z or .65z ?

Nnesha (nnesha):

you*

OpenStudy (howard-wolowitz):

0.65^z

OpenStudy (howard-wolowitz):

why did you say here we go!!!!!!!

Nnesha (nnesha):

"4000⋅0.65z=700 Divide each term in the equation by 4000." okay so here it should be (.65)^z right ?

OpenStudy (howard-wolowitz):

0.65^z=740

OpenStudy (howard-wolowitz):

THEN : Take the natural logarithm of both sides of the equation to remove the variable from the exponent. ln(0.65^z)=ln(740)

Nnesha (nnesha):

and how did you get 740 ?

OpenStudy (howard-wolowitz):

i didnt i got 7/40

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Howard-Wolowitz 0.65^z=740 \(\color{blue}{\text{End of Quote}}\) that's why i asked

OpenStudy (howard-wolowitz):

oh I see .. well thats because I typed it wrong

Nnesha (nnesha):

yes right convert 7/40 into decimal and then take ln both side right!

OpenStudy (howard-wolowitz):

yes

OpenStudy (howard-wolowitz):

sorry I got lost

Nnesha (nnesha):

take ln both side \[\huge\rm ln(.65)^z = \ln(.175)\] and then apply log properties

Nnesha (nnesha):

power rule \[\large\rm ln x^y = y \ln x\] this one

Nnesha (nnesha):

 now divide both side by ln .65 then you would get an answer